# UnrealZoo: Enriching Photo-realistic Virtual Worlds for Embodied AI

### v1

@@ A.1 Comparison with other Simulators
To better explain Table 1, we list the description of each symbol about the scene types and playable entities in Table 6. Since photorealism mainly relies on the engine used, we visualize the snapshots rendered by different engines in Figure 6. Note that Google Maps are images captured in the real world, but can not simulate the dynamic of the scenes and interactions between objects. By utilizing advanced rendering and physics engines, Unreal Engine simulates large-scale photorealistic environments that are not only visually appealing but also capable of complex interactions between agents and objects. So we choose to build environments on Unreal Engine.
@@
Figure 6: Comparison of the visual realism of different engines: we show the snapshots captured from different engines to compare the photo-realism of different environments for an intuitive feeling. Note that Google Maps capture and reconstruct the images from the real world, but can not simulate the dynamic of the scenes and interactions between agents and objects.
@@ A.4 Navigation Mesh
Based on NavMesh, we build an internal navigation system, allowing agents to autonomously navigate with the built-in AI controller in the Unreal Engine. This includes path-finding and obstacle-avoidance capabilities, ensuring smooth and realistic movement throughout diverse terrains and structures. Moreover, in our City style map, we manually construct road segmentation, we manually segment the roads to distinguish between pedestrian and vehicle pathways. When agents use the navigation system for autonomous control, they will navigate the shortest path based on the priority of the different areas. Figure 10 shows an example of the rendered semantic segmentation for NavMesh in an urban city.
@@ B.2 Active Visual Tracking
Referring to previous works (Zhong et al., 2024), we use human characters as an agent player and a continuous action space for agents. The action space contains two variables: the angular velocity and the linear velocity. Angular velocity varies between $-30^{\circ}/s$ and $30^{\circ}/s$, while linear velocity ranges from $-1\ m/s$ to $1\ m/s$. In the agent-centric coordinate system, the reward function is defined as:
@@
where $(\rho,\theta)$ denotes the current target position relative to the tracker, $(\rho^{*},\theta^{*})=(2.5m,0)$ represents the expected target position, i.e., the target should be $2.5m$ in front of the tracker. The error is normalized by the field of the view $(\rho_{max},\theta_{max})$. During execution, an episode ends with a maximum length of 500 steps, applying the appropriate termination conditions. In the experiment, we adopt the original neural network structure and parameters, as listed in Table  9 and  10.
Figure 11: The 8 environments used for collecting offline dataset.
B.3 Task Configuration In a JSON File
We provide an example of the task configuration JSON file in Figure 12. Using the JSON file, we can easily set the configuration of the binary, the continuous and discrete action space for each agent, the placement of the binding camera, choose the area to reset, and other hyper-parameters about the environments.
A Json File for Task Configuration "env_name": env_name, "env_bin":path-to-binary, "env_map": map_name, "env_bin_win": path-to-binary(for windows), "third_cam": {"cam_id": 0,"pitch": -90,"yaw": 0,"roll": 0,"height_top_view": 1460.0,"fov": 90}, "height": 460.0, "interval": 1000, "agents": { "player": { "name": ["BP_Character_923"], "cam_id": [3], "class_name": ["bp_character_C"], "internal_nav": true, "scale": [1,1,1], "relative_location": [20,0,0], "relative_rotation": [ 0,0,0], "head_action_continuous": {"high": [15,15,15], "low": [-15,-15,-15]}, "head_action": [ [0,0,0],[0,30,0],[0,-30,0]], "animation_action": ["stand","jump","crouch"], "move_action": [ [angular, velocity] ... ], "move_action_continuous": {"high": [30,100],"low": [-30,-100]} }, "animal": { "name": ["BP_animal_2"], "cam_id": [1], "class_name": ["BP_animal_C"], "internal_nav": true, "scale": [1,1,1], "relative_location": [20,0,0], "relative_rotation": [0,0,0], "move_action": [ [angular, velocity] ... ], "move_action_continuous": { "high": [30,100],"low": [-30,-100]} }, "drone": { "name": ["BP_Drone01_2"], "cam_id": [2], "class_name": ["BP_drone01_C"], "internal_nav": false, "scale": [ 0.1,0.1,0.1], "relative_location": [0,0,0], "relative_rotation": [0,0,0], "move_action": [ [angular, velocity] ... ], "move_action_continuous": {"high": [1,1,1,1],"low": [-1,-1,-1,-1]} } }, "safe_start": [ [x,y,z], ... ], "reset_area": [x_min,x_maxin,y_min,y_max,z_min,z_max], "random_init": false, "env": {"interactive_door": []}, "obj_num": 466, "size": 192555.0, "area": 9900.0, "bbox": [110.0, 90.0,19.45]
@@ B.3 Task Configuration In a JSON File
B.4 Collecting Demonstration for Active Visual Tracking
To demonstrate the flexibility of the environment, we use state-based expert policy and the multi-level perturbation strategy  (Zhong et al., 2024) to automatically generate various imperfect demonstrations as the offline dataset. For active visual tracking, we employ three distinct datasets for training agents via offline reinforcement learning (Offline RL) algorithms, referred to as 1 Env., 2 Envs., and 8 Envs. The detailed composition of each dataset is depicted in Figure 11. For the 1 Env. dataset, we use only the FlexibleRoom, an abstract environment enriched with diverse augmentation factors, to gather 100k steps of trajectory data. For 2 Envs., we collect 50k step trajectories from FlexibleRoom and an additional 50k steps from the Supermarket environment. The 8 Envs. dataset involves eight different environments, with 12.5k steps collected from each. Therefore, the total amount of data in the three datasets is the same (100k) to ensure the fairness of the comparison. These dataset configurations aim to highlight the critical role of environment diversity in enhancing the generalization capabilities of embodied AI agents.
@@ Appendix C Implementation Details of Agents
C.1 RL-based Agents
Learning to navigate with online reinforcement learning. For navigation, we construct an RL-based end-to-end model, using A3C (Mnih et al., 2016) to accelerate online reinforcement learning in a distributed manner. The model’s structure is as follows: a mask encoder extracts spatial visual features from the segmentation mask, which are then passed to a temporal encoder to capture latent temporal information. Finally, the spatiotemporal features, concatenated with the target’s relative spatial position, are fed into the actor-critic network to optimize the actor layer for action prediction. The detailed network structure and parameters used in the experiment are listed in Table  7 and  8. Here, we provide the training curves in Roof and Factory environments, depicted in Figure 16. In the Factory, we set the number of workers to 4, while in the Roof, the number of workers is set to 6. It can be observed that, for Online RL, the number of workers and the complexity of environments have a significant impact on training efficiency. Looking forward, we anticipate that offline-based algorithms can effectively address the challenges of training efficiency and generalization.
Table 7: Details the neural network structure of RL-based agent for navigation task, where 5$\times$5-32S1 means 32 filters of size 5$\times$5 and stride 1, FC256 indicates the fully connected layer with output dimension 256, and LSTM128 indicates that all the sizes in the LSTM unit are 128.
@@ C.1 RL-based Agents
Parameters 5$\times$5-32S1 2-S2 5$\times$5-32S1 2-S2 4$\times$4-64S1 2-S2 3$\times$3-64S1 2-S2
@@
Table 8: The experiment setting and hyper-parameters used for training the RL-based navigation agent.
@@
Learning to track with offline reinforcement learning. For the tracking task, we adopt an offline reinforcement learning (Offline RL) approach to enhance training efficiency and improve the agent’s generalization to unknown environments. Specifically, we build an end-to-end model trained using offline data and the conservative Q-learning (CQL) strategy (Kumar et al., 2020). We adopt the same model structure from the latest visual tracking agent  (Zhong et al., 2024), consisting of a Mask Encoder, a Temporal Encoder, and an Actor-Critic network. Detailed model structures and training parameters are summarized in Table  9 and  10. Additionally, we provide the model’s loss curves under different dataset setups, as shown in Figure  15. The model achieves near-convergence within two hours across all dataset setups. To ensure the loss curves stabilize fully, we continued training for an additional three hours, during which no significant further decrease in the loss was observed. A comprehensive evaluation of the model’s performance is presented in Tables  11 and  12, highlighting its strong generalization to unseen environments and robustness to dynamic disturbances. The training efficiency, generalization capability, and robustness achieved by offline RL further reinforce our belief that offline RL methods will become a mainstream approach for rapid prototyping and iteration in embodied intelligence systems.
Table 9: Network structure used in the offline RL method (Zhong et al., 2024), where 8$\times$8-16S4 means 16 filters of size 8$\times$8 and stride 4, FC256 indicates a fully connected layer with dimension 256, and LSTM64 indicates that all sizes in the LSTM unit are 64.
@@
Parameters 8$\times$8-16S4 4$\times$4-32S2 256 64 2 2
Table 10: The hyper-parameters used for offline training and the policy network.
Name Value Name Value
Learning Rate 3e-5 LSTM update step 20
Discount Factor 0.99 LSTM Input Dimension 256
Batch Size 32 LSTM Output Dimension 64
LSTM Hidden Layer size 1
C.2 VLM-based Agents
We built agents with a reasoning framework based on the Large Vision-Language Model. We employ OpenAI GPT-4o as the base model. System prompt used in the navigation task, as shown in Figure 14 and system prompt used in the tracking task, as shown in Figure 13.
C.3 Human Benchmark for navigation
@@
System Prompt used for active tracking Objective: You are an intelligent tracking agent designed to control the robot to track the person in the view. The first person in your view is your target. You need to provide concrete moving strategie to helo robot tracking the target in the given environment. Representation details: 1. Moving instructions are concrete actions that the robot can take to adjust its viewpoint and distance to the target. The moving instructions include: -move closer: Move the robot closer to the target. This should be chosen when the target is too far away from the robot and there is no obstacle in the way. -move further: Move the robot further away from the target. This should be 2chosen when the target is too close to the robot and only part of the target body is visible in the view. -keep current: Maintain the current distance and angle between the robot and the target. This is chosen when the target is fully observable in the view and there is enough space in front of both tracker and target without any potential obstacles may cause collision and occlusion. -turn left: Turn the robot to left direction, the target will move towards the right side in next frame. -turn right: Turn the robot to right direction, the target will move towards the left side in next frame. Input Understanding: 1.**Image:** We provide a first-person view observation of the robot to help you understand the surrounding environment. The observation is represented as a color image from the trackers first-person perspective. Output Understanding: 1. **Moving Strategy:** A temporal reasonable move strategy to adjust the robot viewpoint and distance to achieve robotss long-term tracking task. This should be represented as a concrete moving instructions, the instructions should be choose from "move closer", "move further" ,"keep current", "turn left","turn right". Format - [Keep current]. Strategy Considerations: 1.If the persons horizontal position in the robots field of view deviates from the center by more than 25% of the image width, we consider the target to be on one side of the image, otherwise we say the target is near the center. 2.To provide a reasonable moving strategy, you should think step by step based on the input image and the following hints: 1)If the person is too close to the robot and the target in the image is clipped, robot should move further first to obtain a better view. 2)If the persons size in the view is too small in the image, robot should move closer to obtain a better view. 3)If the person may occluded by obstacles or structures in the future, the robot should move closer to avoid losing the person in the next frame. 4)If the person is near the right edge in the image and there is no immediate obstacle in front of robot, the robot should turn right to keep person near center in the image. 5)If there is immediate hinder obstacles in front of the robot, turn right or left to a clean space first. 6)If there is any potential occlusion effect or obstacles on either side of the persons walking path, the robot should move closer to avoid losing the person in the next frame. 7)If there is no person in the current image, turn right or turn left to search the person. Instructions: 1.Provide ONLY the decision in the [output:] strictly following the format without additional explanations or additional text.
@@
System Prompt used for navigation Objective: You are an intelligent navigation agent designed to control the robot to navigate to the target object location based on first-person observation and provide a relative position between the robot and the target. You need to provide an action sequence to help the robot move to the target location. Representation details: 1. Relative Position: This contains three elements, in the format - [Distance, Direction, Height]. -Distance: The relative distance between the robot and the target object. -Direction: The target objects relative direction to the robot, represented in degrees. \ A positive value represtent the target is on the right side of the robot with corresponding angle and a negative value represent the target is on the left side of the robot with corresponding angle. \ The absolute value of the angle larger than 90 degree means the target is behind the robot. \ -Height: The relative vertical position, where a positive value indicates that the target is higher than the robot. 1. Actions: These are the movements the robot can perform to adjust its position. The available actions include: -Move Forward: Propel the robot forward by 100 centimeter. -Move Backward: Propel the robot backward by 100 centimeter. -Turn Left: Rotate the robot 15 degrees to the left. -Turn Right: Rotate the robot 15 degrees to the right. -Jump: Make the robot leap into the air, robot should use this action to jump over obstacles or climb over stairs. -Crouch: Lower the robot into a crouching position for 2 seconds, after which it will automatically stand up. -Keep Current: Maintain the robots current position without any movement. Input Understanding: 1.**Image:** We provide a first-person view observation of the robot to help you understand the surrounding environment. The observation is represented as a color image from the robots first-person perspective. 2.**Relative Position:** This data provides the target objects relative position to the robot, including the distance, direction, and height. The distance is measured in centimeters, the direction in degrees, and the height in centimeters. Output Understanding: 1. **Action Sequence:** This is a series of Three continuous actions that the robot should take to navigate toward the target object. Each sequence must consider the provided relative position data and the first-person observation. \ The actions should be ordered logically to effectively move the robot closer to the target, adjusting its direction, distance, and height as needed. \ The action sequence should be clear and executable, enabling the robot to reach the target efficiently while avoiding obstacles and maintaining stability in the format - [Action1, Action2, Action3]. Each action should be choose from the available actions mentioned above. Strategy Considerations: 1.Assessing Relative Position: Begin by evaluating the target objects relative position in terms of distance, direction, and height to inform the action sequence. 2.Action Combination for Navigation: Utilize the action sequence to create effective combinations, each action will last for 1 seconds. For example: -Consider using multiple consecutive actions like [Move Forward, Jump, Jump] to climb over the front obstacles or boxes. -Consider using [Move Backward,Move Backward,Move Backward] to move the robot avoid a front wall or fence. 3.Obstacle Detection: Leverage the first-person observation to identify obstacles. Based on their location, formulate action sequences that facilitate smooth navigation while avoiding collisions. 4.Efficient Pathing: Ensure the action sequence is designed to dynamically adjust the robot movement torward target object, which is minimize the distance and direction value in **Relative Position**. 5.Sequence Validation: Validate the generated action sequence and consider past memories to ensure it is practical given the current environment and obstacles, making long-term adjustments as necessary. Instructions: 1.Provide ONLY the action sequence in the [output:] strictly following the format -[Action1, Action2, Action3], without additional explanations or additional text.

### v2

@@ A.1 Comparison with other Simulators
To better explain Table 1, we list the description of each symbol about the scene types and playable entities in Table 8. Since photorealism mainly relies on the engine used, we visualize the snapshots rendered by different engines in Figure 7. Note that Google Maps are images captured in the real world, but can not simulate the dynamics of the scenes and interactions between objects. By utilizing advanced rendering and physics engines, Unreal Engine simulates large-scale photorealistic environments that are not only visually appealing but also capable of complex interactions between agents and objects. So we choose to build environments on Unreal Engine.
Table 6: We evaluate the simulation speed on an Nvidia GTX 4090 GPU, Intel i7-14700K CPU, and Windows OS by repeatedly calling UnrealCV+ API function 1000 times and recording the execution time. Multi-agent interaction performance is measured using the gym interface. The image resolution is set to 640×480. The table presents the FPS measurement results across environments of varying scales.
Env Color Image Object Mask Surface Normal Depth Image 2 Agents Interaction 6 Agents Interaction 10 Agents Interaction
FlexibleRoom (71 objects, 2440m2) 85 164 137 100 94 40 27
BrassGarden (467 objects, 9900m2) 107 214 173 123 102 48 31
Supermarket (2839 objects, 11700m2) 99 173 167 117 70 33 19
SuburbNeighborhoodDay (2469 objects, 23100m2) 79 139 112 92 53 27 17
GreekIsland (3174 objects, 448800m2) 82 167 124 103 52 23 16
MedievalNatureEnvironment (8534 objects, 16km2) 70 112 97 74 29 14 10
Table 7: The comparison with related photo-realistic virtual worlds for embodied AI.
@@
Figure 7: Comparison of the visual realism of different engines: we show the snapshots captured from different engines to compare the photo-realism of different environments for an intuitive feeling. Note that Google Maps captures and reconstructs the images from the real world, but can not simulate the dynamics of the scenes and interactions between agents and objects.
@@ A.4 Navigation Mesh
Based on NavMesh, we build an internal navigation system, allowing agents to autonomously navigate with the built-in AI controller in the Unreal Engine. This includes path-finding and obstacle-avoidance capabilities, ensuring smooth and realistic movement throughout diverse terrains and structures. Moreover, in our City style map, we manually construct road segmentation, and we manually segment the roads to distinguish between pedestrian and vehicle pathways. When agents use the navigation system for autonomous control, they will navigate the shortest path based on the priority of the different areas. Figure 10 shows an example of the rendered semantic segmentation for NavMesh in an urban city.
@@ B.2 Active Visual Tracking
Referring to previous works [49], we use human characters as an agent player and a continuous action space for agents. The action space contains two variables: the angular velocity and the linear velocity. Angular velocity varies between $30^{\circ}/s$ and $30^{\circ}/s$, while linear velocity ranges from 1m/s-1\ m/s to 1m/s1\ m/s. In the agent-centric coordinate system, the reward function is defined as:
@@
where $(\rho,\theta)$ denotes the current target position relative to the tracker, $(\rho^{*},\theta^{*})=(2.5m,0)$ represents the expected target position, i.e., the target should be $2.5m$ in front of the tracker. The error is normalized by the field of view $(\rho_{max},\theta_{max})$. During execution, an episode ends with a maximum length of 500 steps, applying the appropriate termination conditions. In the experiment, we adopt the original neural network structure and parameters, as listed in Table  11 and  12.
Figure 11: The 8 environments used for collecting the offline dataset.
B.3 Task Configuration in JSON File
We provide an example of the task configuration JSON file in Figure 12. Using the JSON file, we can easily set the configuration of the binary, the continuous and discrete action space for each agent, the placement of the binding camera, choose the area to reset, and other hyperparameters about the environments.
A Json File for Task Configuration env_name”: env_name, env_bin”:path-to-binary, env_map”: map_name, env_bin_win”: path-to-binary(for windows), third_cam”: {”cam_id”: 0,”pitch”: -90,”yaw”:0,”roll”:0,”height_top_view”: 1460.0,”fov”: 90}, height”: 460.0, interval”: 1000, agents”: { player”: { name”: [”BP_Character_923”], cam_id”: [3], class_name”: [”bp_character_C”], internal_nav”: true, scale”: [1,1,1], relative_location”: [20,0,0], relative_rotation”: [ 0,0,0], head_action_continuous”: {”high”: [15,15,15], low”: [-15,-15,-15]}, head_action”: [ [0,0,0],[0,30,0],[0,-30,0]], animation_action”: [”stand”,”jump”,”crouch”], move_action”: [ [angular, velocity] ], move_action_continuous”: {”high”: [30,100],”low”: [-30,-100]} }, animal””{ } drone”: { } }, safe_start”: [ [x,y,z], ], reset_area”: [x_min,x_maxin,y_min,y_max,z_min,z_max], random_init”: false, env”: {”interactive_door”: []}, obj_num”: 466, size”: 192555.0, area”: 9900.0, bbox”: [110.0, 90.0,19.45]
@@ Appendix C Implementation Details of Agents
C.1 Data Collection for Offline RL
To collect demonstrations for offline reinforcement learning, we use a state-based expert policy and the multi-level perturbation strategy  [49] to automatically generate various imperfect demonstrations as the offline dataset. For active visual tracking, we employ three distinct datasets for training agents via offline reinforcement learning (Offline RL) algorithms, referred to as 1 Env., 2 Envs., and 8 Envs. The detailed composition of each dataset is depicted in Figure 11. For the 1 Env. dataset, we use only the FlexibleRoom, an abstract environment enriched with diverse augmentation factors, to gather 100k steps of trajectory data. For 2 Envs., we collect 50k step trajectories from FlexibleRoom and an additional 50k steps from the Supermarket environment. The 8 Envs. dataset involves eight different environments, with 12.5k steps collected from each. Therefore, the total amount of data in the three datasets is the same (100k) to ensure the fairness of the comparison. These dataset configurations aim to highlight the critical role of environment diversity in enhancing the generalization capabilities of embodied AI agents.
C.2 RL-based Agents
Learning to navigate with online reinforcement learning. For navigation, we construct an RL-based end-to-end model, using A3C [27] to accelerate online reinforcement learning in a distributed manner. The model’s structure is as follows: a mask encoder extracts spatial visual features from the segmentation mask, which are then passed to a temporal encoder to capture latent temporal information. Finally, the spatiotemporal features, concatenated with the target’s relative spatial position, are fed into the actor-critic network to optimize the actor layer for action prediction. The detailed network structure and parameters used in the experiment are listed in Table  9 and  10. Here, we provide the training curves in Roof and Factory environments, depicted in Figure 16. In the Factory, we set the number of workers to 4, while in the Roof, the number of workers is set to 6. It can be observed that, for Online RL, the number of workers and the complexity of environments have a significant impact on training efficiency. Looking forward, we anticipate that offline-based algorithms can effectively address the challenges of training efficiency and generalization.
Table 9: Details the neural network structure of the RL-based agent for the navigation task, where $\times5$-32S1 means 32 filters of size $\times5$ and stride 1, FC256 indicates the fully connected layer with output dimension 256, and LSTM128 indicates that all the sizes in the LSTM unit are 128.
@@ C.2 RL-based Agents
Parameters $\times5$-32S1 2-S2 $\times5$-32S1 2-S2 $\times4$-64S1 2-S2 $\times3$-64S1 2-S2
@@
Table 10: The experiment setting and hyperparameters used for training the RL-based navigation agent.
@@
Learning to track with offline reinforcement learning. For the tracking task, we adopt an offline reinforcement learning (Offline RL) approach to enhance training efficiency and improve the agent’s generalization to unknown environments. Specifically, we build an end-to-end model trained using offline data and the conservative Q-learning (CQL) strategy [21]. We adopt the same model structure from the latest visual tracking agent  [49], consisting of a Mask Encoder, a Temporal Encoder, and an Actor-Critic network. Detailed model structures and training parameters are summarized in Table 11 and  12. Additionally, we provide the model’s loss curves under different dataset setups, as shown in Figure  15. The model achieves near-convergence within two hours across all dataset setups. To ensure the loss curves stabilize fully, we continued training for an additional three hours, during which no significant further decrease in the loss was observed. A comprehensive evaluation of the model’s performance is presented in Tables  13 and  14, highlighting its strong generalization to unseen environments and robustness to dynamic disturbances. The training efficiency, generalization capability, and robustness achieved by offline RL further reinforce our belief that offline RL methods will become a mainstream approach for rapid prototyping and iteration in embodied intelligence systems.
Table 11: Network structure used in the offline RL method [49], where $\times8$-16S4 means 16 filters of size $\times8$ and stride 4, FC256 indicates a fully connected layer with dimension 256, and LSTM64 indicates that all sizes in the LSTM unit are 64.
@@
Parameters $\times8$-16S4 $\times4$-32S2 256 64 2 2
Table 12: The hyperparameters used for offline training and the policy network.
Parameter Value Parameter Value
Learning Rate 3e-5 LSTM steps 20
Discount 0.99 LSTM In Dim 256
Batch Size 32 LSTM Out Dim 64
LSTM Hidden Layers 1
C.3 VLM-based Agents
We built agents with a reasoning framework based on the Large Vision-Language Model. We employ OpenAI GPT-4o as the base model. System prompt used in the navigation task, as shown in Figure 14, and system prompt used in the tracking task, as shown in Figure 13.
C.4 Human Benchmark for Navigation
@@
[…]
@@
System Prompt used for navigation Objective: You are an intelligent navigation agent designed to control the robot to navigate to the target object location based on first-person observation and provide a relative position between the robot and the target. You need to provide an action sequence to help the robot move to the target location. Representation details: 1. Relative Position: This contains three elements, in the format - [Distance, Direction, Height]. -Distance: The relative distance between the robot and the target object. -Direction: The target objects relative direction to the robot, represented in degrees. 1. Actions: These are the movements the robot can perform to adjust its position. The available actions include: -Move Forward: Propel the robot forward by 100 centimeter. -Move Backward: Propel the robot backward by 100 centimeter. -Turn Left: Rotate the robot 15 degrees to the left. -Turn Right: Rotate the robot 15 degrees to the right. -Jump: Make the robot leap into the air, robot should use this action to jump over obstacles or climb over stairs. -Crouch: Lower the robot into a crouching position for 2 seconds, after which it will automatically stand up. -Keep Current: Maintain the robots current position without any movement. Input Understanding: 1.**Image:** We provide a first-person view observation of the robot to help you understand the surrounding environment. The observation is represented as a color image from the robots first-person perspective. 2.**Relative Position:** This data provides the target objects relative position to the robot, including the distance, direction, and height. The distance is measured in centimeters, the direction in degrees, and the height in centimeters. Output Understanding: 1. **Action Sequence:** This is a series of Three continuous actions that the robot should take to navigate toward the target object, in the format - [Action1, Action2, Action3]. Each action should be chosen from the available actions mentioned above. Strategy Considerations: 1.Assessing Relative Position: Begin by evaluating the target objects relative position in terms of distance, direction, and height to inform the action sequence. 2.Action Combination for Navigation: Utilize the action sequence to create effective combinations, each action will last for 1 seconds. 3.Obstacle Detection: Leverage the first-person observation to identify obstacles. Based on their location, formulate action sequences that facilitate smooth navigation while avoiding collisions. Instructions: 1.Provide ONLY the action sequence in the [output:] strictly following the format -[Action1, Action2, Action3], without additional explanations or additional text.